OpenStudy (anonymous):

radical equations help!! sqrt x+2+4=x the square root is over the x+2 i got -7 and -2 but i plugged them in and i dont know where i went wrong but they didnt work

OpenStudy (jdoe0001):

\(\bf \sqrt{x+2}+4=x\sqrt{x+2}\ \ \ ?\)

OpenStudy (anonymous):

why would you do that?

OpenStudy (wolf1728):

I wrote a graphic and the solutions I got were X=+7 and X=+2